OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (amistre64):

you might need to go over your 5 times tables for this one ....

OpenStudy (anonymous):

is there nothing special i do for this type of case?

OpenStudy (amistre64):

the only thing you need to do is follow the rule they provide

OpenStudy (amistre64):

a1 = -8 and each new term is 5 times the previous term

OpenStudy (anonymous):

oh der...

OpenStudy (amistre64):

a1 = -8 a2 = 5 * a1 a3 = 5 * a2 a4 = 5 * a3 a5 = 5 * a4 a6 = 5 * a5

OpenStudy (anonymous):

thank you!!! I feel dumb now lol

OpenStudy (amistre64):

:)